OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. I need help with the square root of the equation. I will put the equation down below.

OpenStudy (anonymous):

\[f(x)=x-3\]\[g(x)=\sqrt{x+3}\]

OpenStudy (anonymous):

sorry it's\[f(x)=x^2-3\]